Bubba:

Compare 8 ⋅ 107 to 2 ⋅ 104. 8 ⋅ 107 is 4,000 times larger than 2 ⋅ 104. 8 ⋅ 107 is 400 times larger than 2 ⋅ 104. 8 ⋅ 107 is 40 times larger than 2 ⋅ 104. 8 ⋅ 107 is 4 times larger than 2 ⋅ 104.

Bubba:

@volpina @Vocaloid the bottom one???

Vocaloid:

Hint: when dividing powers of the same base, subtract exponents (10^7)/10^4 =

Bubba:

7-4=3

Bubba:

so now what?

Vocaloid:

Yes so 10^7/10^4 = 10^3 So your answer is 4 * 10^3 = ?

Bubba:

the one above the bottom?

Bubba:

i mean the top one?

Bubba:

@Vocaloid ? Top one?

Vocaloid:

Yes good

Vocaloid:

I'm on my phone run so it's a little slow

Bubba:

okay I have 4 more comming